# Grounded-VideoLLM: Sharpening Fine-grained Temporal Grounding in Video Large Language Models

### v1

@@ Abstract
Video Large Language Models (Video-LLMs) have demonstrated remarkable capabilities in coarse-grained video understanding, however, they struggle with fine-grained temporal grounding. In this paper, we introduce Grounded-VideoLLM, a novel Video-LLM adept at perceiving and reasoning over specific video moments in a fine-grained manner. We identify that current Video-LLMs have limitations for fine-grained video understanding since they lack effective temporal modeling and timestamp representation. In light of this, we sharpen our model by incorporating (1) an additional temporal stream to encode the relationships between frames and (2) discrete temporal tokens enriched with specific time knowledge to represent timestamps. To optimize the training of Grounded-VideoLLM, we employ a multi-stage training scheme, beginning with simple video-captioning tasks and progressively introducing video temporal grounding tasks of increasing complexity. To further enhance Grounded-VideoLLM’s temporal reasoning capability, we also curate a grounded VideoQA dataset by an automatic annotation pipeline. Extensive experiments demonstrate that Grounded-VideoLLM not only excels in fine-grained grounding tasks such as temporal sentence grounding, dense video captioning, and grounded VideoQA, but also shows great potential as a versatile video assistant for general video understanding.
@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) have made remarkable progress in image-level understanding (Liu et al., 2023; Dai et al., 2023; Li et al., 2023a). However, extending their capabilities to the video domain poses distinct challenges. Unlike static images, the temporal nature of videos challenges models to process not only visual content but also the sequence and timing of events. While current Video-LLMs (Xu et al., 2024a; Li et al., 2024; Zhang et al., 2023b; Lin et al., 2023) are capable of capturing global visual semantics and generating coarse-grained captions for short clips, they struggle with fine-grained video understanding (Liu et al., 2024b; Wang et al., 2024d), which requires decomposing the video along the temporal axis to accurately perceive and reason over specific moments, such as subtle actions, transitions, and events that unfold over time.
@@
Figure 1: Grounded-VideoLLM enables Temporal Referring/Localizing/Reasoning for MLLM.
@@
Previous research efforts (Ren et al., 2024; Huang et al., 2024a; Qian et al., 2024a; Huang et al., 2024b; Guo et al., 2024) have explored temporal grounding to improve fine-grained video understanding. However, two main challenges impede their potential for achieving effective fine-grained temporal grounding: (1) Models like Video-ChatGPT (Maaz et al., 2024b), P-LLaVA (Xu et al., 2024a), and Video-LLAMA (Zhang et al., 2023b) typically sample multiple frames from a video and encode each frame independently using an image encoder, followed by a feature projector (e.g., sliding Q-former (Ren et al., 2024), slot-based token compression (Guo et al., 2024), or visual adapter (Huang et al., 2024a)). This approach focuses primarily on spatial details while potentially neglecting the temporal relationships between frames (Maaz et al., 2024a). (2) Current models also struggle with timestamp representation, which is crucial for pinpointing specific moments in time for fine-grained understanding. Models such as TimeChat (Ren et al., 2024) and VTimeLLM (Huang et al., 2024a) represent timestamps as plain texts, for example, ["from 102.3 to 120.1 seconds"]. Despite being straightforward, this strategy needs to tokenize continuous floating-point values, which is inefficient for LLMs since their next-token prediction paradigm struggles with handling numerical data (Schwartz et al., 2024; Frieder et al., 2023).
To further improve video comprehension, we propose to sharpen the model with fine-grained temporal grounding, which allows the model to not only recognize what happens but also pinpoint when it happens in a fine-grained manner. We identify three essential capabilities of fine-grained temporal grounding, as illustrated in Figure 1, including Temporal Referring (Huang et al., 2024a), Localizing (Gao et al., 2017; Caba Heilbron et al., 2015), and Reasoning (Xiao et al., 2024; Qian et al., 2024a). Referring involves the model precisely describing events within user-specified time intervals. Localizing challenges the model to identify the timestamps of a given query and construct a coherent storyline by locating a sequence of events within the video. Reasoning combines both Referring and Localizing capabilities and further requires the ability to answer questions by understanding temporal dynamics and applying wrold knowledge. Achieving these tasks necessitates both effective temporal modeling and robust timestamp representation.
@@
Targeting these goals, we introduce Grounded-VideoLLM, a novel Video-LLM that can perceive and reason over specific video moments with fine-grained precision. From the perspective of model architecture, Grounded-VideoLLM is built upon two key innovations: (1) Two-Stream Encoding: We decompose each segment of the video into spatial and temporal components, and encode each with an expert encoder respectively. The temporal stream extracts motion representations from dense frames and complements the spatial stream which captures appearance representations. This dual-stream approach forms comprehensive video representations enriched with both temporal and spatial information. (2) Temporal Tokens: We extend the LLM’s vocabulary by introducing discrete tokens specifically crafted for timestamp representation. These temporal tokens denote relative time positions within a video and share a unified embedding space with the LLM. This integration allows Grounded-VideoLLM to avoid the inefficiency of tokenizing numerical text, enabling seamless prediction of both timestamps and textual outputs in a single sequence of discrete tokens. From the perspective of training and datasets, given that such a unified model requires extensive pre-training, we start with an image-based MLLM (Microsoft, 2024) as the foundation and adopt a three-stage training strategy to ensure efficiency. We meticulously select different tasks for each stage, and progressively refine the model in a “coarse-to-fine” manner, transitioning from image understanding to video comprehension, and ultimately to fine-grained temporal grounding. Furthermore, we enhance the model’s temporal reasoning capability by curating 17K grounded VideoQA (Xiao et al., 2024) samples with the assistance of GPT-4. Extensive experiments demonstrate that Grounded-VideoLLM shows promising results over existing Video-LLMs not only in traditional video temporal grounding tasks but also in general video understanding benchmarks.
In summary, we make the following contributions: (1) We present key insights into fine-grained temporal grounding for Video-LLMs and introduce Grounded-VideoLLM, a Video-LLM adept at perceiving and reasoning over specific video moments. This is achieved through a two-stream architecture for effective temporal modeling and the temporal tokens for efficient timestamp representation. (2) We propose a step-by-step training strategy that progressively adapts an image-based MLLM into a robust Video-LLM, while curating a grounded VideoQA dataset to further enhance the temporal reasoning capability. (3) We conduct extensive experiments across various tasks, including temporal sentence grounding, dense video caption, VideoQA, and general Video-LLM benchmarks, demonstrating the superiority of Grounded-VideoLLM for fine-grained video understanding.
@@ 2 Related Work
Video Large Language Models have caught a growing interest with the advancements in image-based MLLMs (Zhang et al., 2023b; Lin et al., 2023; Maaz et al., 2024a; Luo et al., 2023). Despite promising results, current Video-LLMs, such as P-LLaVA (Xu et al., 2024a) and Video-ChatGPT (Maaz et al., 2024b), often struggle with temporal understanding (Liu et al., 2024b) and exhibit temporal hallucination (Wang et al., 2024d) when answering questions about specific video moments. These models encode each frame independently using a pre-trained image encoder, concatenating the frame embeddings to create a video representation. This late fusion method can result in video representations that lack inherent temporal information and heavily rely on the position embeddings of LLM (Su et al., 2024) for temporal understanding, limiting the model’s capability to perform fine-grained temporal grounding and understanding. In contrast to these studies, we employ a two-stream architecture that integrates a video expert to extract motion features to complement the appearance features during the early encoding process. Additionally, we employ a progressive training strategy that gradually adapts an image-based MLLM for fine-grained video understanding. Unlike concurrent studies such as SlowFast-LLaVA (Xu et al., 2024b) and VideoGPT+ (Maaz et al., 2024a), which also utilize a two-stream architecture, we specifically targets fine-grained temporal grounding through a unique encoding/pooling/training strategy for dense frames and grounding design.
@@
Video Temporal Grounding (VTG) aims to associate specific video moments with their corresponding timestamps. Traditional VTG tasks include Temporal Sentence Grounding (Gao et al., 2017; Hendricks et al., 2018) and Dense Video Captioning (Caba Heilbron et al., 2015; Zhou et al., 2018). Other tasks, such as Grounded VideoQA (Xiao et al., 2024), emphasize reasoning over videos with grounded information. Given the emerging reasoning capabilities of Video-LLMs, many studies have investigated how to adapt them for VTG tasks. For example, TimeChat (Ren et al., 2024), VTimeLLM (Huang et al., 2024a), and LITA (Huang et al., 2024b) perform temporal grounding using a fully text-to-text approach through instruction-tuning datasets. Momentor (Qian et al., 2024a) introduces a temporal perception module to address the quantization errors associated with time tokens, while VTG-LLM (Guo et al., 2024) incorporates a limited set of absolute-time tokens to handle timestamp knowledge. Compared to these studies, we avoid textual representation of timestamps and instead introduce discrete and relative temporal tokens for more efficient timestamp encoding. While Vid2Seq (Yang et al., 2023) also employs specialized tokens to indicate temporal positions, it relies heavily on large-scale pre-training from scratch using noisy transcribed speech and is limited to dense video captioning. Our two-stream architecture and progressive training strategy, however, enable MLLMs to efficiently comprehend videos and effectively handle diverse tasks of fine-grained temporal referring, localizing, and reasoning.
@@ 3.1 Two-Stream Encoding
To effectively model the long-range temporal structure, given a video $\mathcal{V}$ with $T$ frames, we divide it into $K$ segments and employ a segment-wise encoding strategy (Wang et al., 2016). Due to the inherent redundancy of consecutive frames in videos, each segment can be naturally represented from two perspectives: spatial and temporal. The spatial representation of each segment is derived from an individual keyframe, capturing the primary appearance semantics, while the temporal representation is learned from multiple frames depicting the motion evolution within the segment. We discuss the details of our two-stream segment encoding as follows.
@@
Spatial Stream. We sample the middle frame from each segment as the keyframe and extract its spatial features using the original image encoder from the MLLM (Radford et al., 2021), resulting in spatial features $\textbf{F}_{S}\in\mathbb{R}^{H_{S}\times W_{S}\times D_{S}}$, where $H_{S}$, $W_{S}$, $D_{S}$ denote the height, width and dimension of the spatial features. Since dense frames are crucial for fine-grained temporal grounding, an appropriate pooling strategy is required to reduce token length for efficient computation. As indicated by Xu et al. (2024a) and Yao et al. (2024) that a 2D average pooling is both efficient and robust for spatial downsampling, we employ a 2D pooling kernel with a size $\sigma_{S}\times\sigma_{S}$ over the feature map and gets $\textbf{F}_{S}\in\mathbb{R}^{N_{S}\times D_{S}}$ as the feature for spatial stream, where $N_{S}=\frac{H_{S}}{\sigma_{S}}\times\frac{W_{S}}{\sigma_{S}}$.
@@
Temporal Stream. Traditional two-stream networks (Simonyan & Zisserman, 2014; Feichtenhofer et al., 2016) typically encode the optical flow as the temporal stream. However, given the scale of data and parameters involved in MLLMs, extracting optical flow is computationally expensive and impractical. Consequently, we resort to a strong and well pre-trained video encoder, InternVideo2 (Wang et al., 2024b), to extract motion representations for each segment, using a lower resolution but more frames. Specifically, we input each segment, containing $\frac{T}{K}$ frames, into the video encoder to obtain the segment-level features $\textbf{F}_{T}\in\mathbb{R}^{\frac{T}{K}\times H_{T}\times W_{T}\times D_{T}}$, where $H_{T}$, $W_{T}$, $D_{T}$ denote the height, width and dimension of each frame feature. Similar to the spatial stream, we apply a 2D average pooling strategy to downsample $\textbf{F}_{T}$. However, as the temporal stream focuses primarily on temporal modeling, we retain the complete temporal information by only pooling along the spatial dimensions. Specifically, we aggressively downsample $\textbf{F}_{T}$ using a kernel with a larger size of $\sigma_{T}\times\sigma_{T}$, resulting in the compressed $\textbf{F}_{T}\in\mathbb{R}^{\frac{T}{K}\times N_{T}\times D_{T}}$ for temporal stream, where $N_{T}=\frac{H_{T}}{\sigma_{T}}\times\frac{W_{T}}{\sigma_{T}}$.
To get the complete segment-level representation, we flatten the features of the spatial stream and temporal stream and concat them together :
$\textbf{F}_{Seg}=\operatorname{Concat}\left[\operatorname{Flatten}(f(\textbf{F}_{S}));\operatorname{Flatten}(g(\textbf{F}_{T}))\right],\textbf{F}_{Seg}\in\mathbb{R}^{(N_{S}+\frac{T}{K}\cdot N_{T})\times D}$ (1)
@@
where $f(\cdot)$ and $g(\cdot)$ are two MLPs that project the visual features to LLM’s dimension $D$. The final video representation is formed by concatenating the $K$ segment-level representations $\textbf{F}_{Seg}$, resulting in $\textbf{F}_{Vid}\in\mathbb{R}^{K\cdot(N_{S}+\frac{T}{K}\cdot N_{T})\times D}$. This representation retains detailed spatial information across all segments along with their global temporal contexts, while maintaining a manageable token length. The combined video representation $\textbf{F}_{Vid}$ is then fed into the LLM serving as soft prompts, alongside the word embeddings of the instruction text $\textbf{F}_{Text}$ to generate the target response $\mathcal{A}$. The model is trained using the cross-entropy loss function with trainable parameters $\theta$:

### v2

@@ Abstract
Despite their impressive performance in coarse-grained video understanding, Video Large Language Models (Video-LLMs) still face challenges in fine-grained temporal grounding, including ineffective temporal modeling and inadequate timestamp representations. In this work, we introduce Grounded-VideoLLM, a novel Video-LLM designed to perceive and reason over specific video moments with fine-grained temporal precision. Our model features (1) a two-stream encoder that explicitly captures inter-frame relationships while preserving intra-frame visual details and (2) discrete temporal tokens enriched with structured time knowledge for timestamp representation. Besides, we propose a multi-stage training strategy tailored to such grounding-specific architecture. The model is initially trained on simple video-caption tasks and progressively introduced to complex video temporal grounding tasks, ensuring a smooth learning curve and temporal alignment. We further strengthen Grounded-VideoLLM’s temporal reasoning by constructing a VideoQA dataset with grounded information using an automated annotation pipeline. Extensive experiments demonstrate that Grounded-VideoLLM not only surpasses existing models in fine-grained grounding tasks but also exhibits strong potential as a general video understanding assistant.
[…]
Grounded-VideoLLM: Sharpening Fine-grained Temporal Grounding in Video Large Language Models
Haibo Wang1,6thanks: Work done during an internship at UC Davis.,  Zhiyang Xu2, Yu Cheng3, Shizhe Diao4, Yufan Zhou5, Yixin Cao6, Qifan Wang7, Weifeng Ge6, Lifu Huang1 thanks: Corresponding author. Codes will be available at https://github.com/WHB139426/Grounded-Video-LLM 1University of California, Davis  2Virginia Tech  3The Chinese University of Hong Kong 4NVIDIA  5Adobe Research  6Fudan University  7Meta AI hibwang@ucdavis.edu, lfuhuang@ucdavis.edu
@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) have made remarkable progress in image-level understanding (Liu et al., 2023; Dai et al., 2023; Li et al., 2023b). However, extending their capabilities to the video domain poses distinct challenges. Unlike static images, the temporal nature of videos challenges models to process not only visual content but also the sequence and timing of events. While current Video-LLMs (Xu et al., 2024a; Li et al., 2024; Zhang et al., 2023b; Lin et al., 2023) are capable of capturing global visual semantics and generating coarse-grained captions for short clips, they struggle with fine-grained video understanding (Liu et al., 2024c; Wang et al., 2024d), which requires decomposing the video along the temporal axis to accurately perceive and reason over specific moments, such as subtle actions, transitions, and events that unfold over time.
@@
Figure 1: Grounded-VideoLLM enables Temporal Referring/Localizing/Reasoning for MLLMs.
@@
Previous research efforts (Ren et al., 2024; Huang et al., 2024a; Qian et al., 2024a; Huang et al., 2024b; Guo et al., 2024) have explored temporal grounding to improve fine-grained video understanding. However, two main challenges impede their potential for achieving effective fine-grained temporal grounding: (1) Models like Video-ChatGPT (Maaz et al., 2024b), P-LLaVA (Xu et al., 2024a), and Video-LLAMA (Zhang et al., 2023b) typically sample multiple frames from a video and encode each frame independently using an image encoder, followed by a feature projector (e.g., sliding Q-former (Ren et al., 2024), visual adapter (Huang et al., 2024a)). This focuses primarily on spatial details while potentially neglecting the temporal relationships between frames since their visual encoders are solely trained on images. (2) Current models also struggle with timestamp representation, which is crucial for pinpointing specific moments in time for fine-grained understanding. Models such as TimeChat (Ren et al., 2024) and VTimeLLM (Huang et al., 2024a) represent timestamps as plain texts, for example, ["from 102.3 to 120.1 seconds"]. Despite being straightforward, this needs to tokenize continuous floating-point values, which is inefficient for LLMs since their next-token prediction paradigm struggles with handling numerical data (Schwartz et al., 2024; Frieder et al., 2023). Although there have been some previous works (Yang et al., 2023; Huang et al., 2024b; Qian et al., 2024a) using special tokens to represent time positions, Vid2Seq Yang et al. (2023) relies heavily on large-scale pre-training from scratch using noisy transcribed speech and is limited to the task of dense video captioning, while LITA Huang et al. (2024b) and Momentor Qian et al. (2024a) only align these tokens with simple fine-tuning stage, which proves to be insufficient in our experiments (Table 5).
@@
To further improve video understanding, we propose to sharpen the model with fine-grained temporal grounding, allowing the model to recognize not only what happens but pinpoint when it happens with finer granularity, as illustrated in Figure 1. Targeting these goals, we introduce Grounded-VideoLLM, a novel Video-LLM that can perceive and reason over specific video moments with fine-grained precision. From the perspective of model architecture, Grounded-VideoLLM is built upon two key innovations: (1) Two-Stream Encoding: We decompose each segment of the video into spatial and temporal components and encode each with an expert encoder. The temporal stream extracts motion representations from dense frames and complements the spatial stream, which captures appearance representations. This dual-stream approach forms comprehensive video representations enriched with both temporal and spatial information. (2) Temporal Tokens: We extend the LLM’s vocabulary by introducing discrete tokens crafted to denote relative time positions and share a unified embedding space with the LLM, allowing Grounded-VideoLLM to avoid the inefficiency of tokenizing numerical text and seamlessly predict both timestamps and textual outputs in a single sequence of discrete tokens. From the perspective of training, we start with an image MLLM (Microsoft, 2024) as the foundation and adopt a three-stage training strategy. We meticulously select different tasks for each stage, and progressively refine the model in a “coarse-to-fine” manner, transitioning from image understanding to video comprehension, and ultimately to fine-grained temporal grounding. This scheme ensures the introduced temporal tokens align closely with the video timelines and LLM’s semantic space, distinguishing our method from previous studies Huang et al. (2024b); Qian et al. (2024a); Yang et al. (2023). Furthermore, we enhance the model’s temporal reasoning by curating 17K grounded VideoQA (Xiao et al., 2024) samples with the assistance of GPT-4 Achiam et al. (2023). Extensive experiments demonstrate that Grounded-VideoLLM shows promising results over existing Video-LLMs not only in traditional video temporal grounding tasks but also in general video understanding benchmarks.
@@ 2 Related Work
Video Large Language Models have caught a growing interest for general video understanding (Zhang et al., 2023b; Lin et al., 2023). However, they struggle with temporal perception (Liu et al., 2024c) and exhibit hallucination (Wang et al., 2024d) when asked about specific moments. They typically encode each frame independently using an image encoder to create video embeddings, resulting in video representations lacking inherent temporal information and heavily relying on the position embeddings of LLM for temporal understanding, limiting the model’s capability to perform fine-grained temporal grounding. Norton (Lin et al., 2024) employs video-paragraph and clip-caption contrastive losses to capture long-term dependencies for fine-grained alignment. In this work, we employ a two-stream architecture that integrates a video expert to extract motion features to complement the appearance features during the early encoding process. This is different from traditional two-stream networks (Simonyan and Zisserman, 2014; Feichtenhofer et al., 2016) since we don’t rely on heavy extraction of optical flows. Additionally, we employ a progressive training strategy that gradually adapts an image-based MLLM for fine-grained video understanding. Although concurrent works such as SlowFast-LLaVA (Xu et al., 2024b) and VideoGPT+ (Maaz et al., 2024a) also introduce an additional stream, SlowFast-LLaVA relies on a single image encoder to process each video frame without training, missing crucial temporal relationships between frames. VideoGPT+ merely arranges video tokens as a prefix to image tokens using sparse frames. Instead, our approach is specifically designed for fine-grained temporal grounding, leveraging a unique encoding, pooling, and training strategy tailored for dense frames, along with a dedicated grounding mechanism.
@@
Video Temporal Grounding (VTG) tasks usually include Temporal Sentence Grounding (Gao et al., 2017; Hendricks et al., 2018), Dense Video Captioning (Caba Heilbron et al., 2015; Zhou et al., 2018), and Grounded VideoQA (Xiao et al., 2024). Given the emerging capabilities of Video-LLMs (Zhang et al., 2023b; Lin et al., 2023; Wang et al., 2024b), many studies have investigated how to adapt them for VTG tasks. For example, TimeChat (Ren et al., 2024) and VTimeLLM (Huang et al., 2024a) perform temporal grounding using a fully text-to-text approach through instruction-tuning datasets. Momentor (Qian et al., 2024a) introduces a temporal perception module to address the quantization errors, and VTG-LLM (Guo et al., 2024) incorporates absolute-time tokens to handle timestamps. Compared to these, we avoid textual representation of timestamps and instead introduce discrete temporal tokens for timestamp encoding. Different from previous methods that also use special tokens Huang et al. (2024b); Qian et al. (2024a); Yang et al. (2023); Peng et al. (2023), our model is more efficient by continuing training based on an established image MLLM with a two-stream architecture and a progressive training strategy.
@@ 3.1 Two-Stream Encoding
Given a video $\mathcal{V}$ with $T$ frames, we divide it into $K$ segments and employ a segment-wise encoding strategy. Due to the inherent redundancy of consecutive frames, each segment can be naturally represented from two perspectives: spatial and temporal. The spatial representation of each segment is derived from an individual keyframe, capturing the primary appearance semantics, while the temporal representation is learned from multiple frames depicting the motion evolution within the segment.
@@
Spatial Stream. We sample the middle frame from each segment as the keyframe and extract its spatial features using the original image encoder from the MLLM (Radford et al., 2021), resulting in spatial features $\textbf{F}_{S}\in\mathbb{R}^{H_{S}\times W_{S}\times D_{S}}$, where $H_{S}$, $W_{S}$, $D_{S}$ denote the height, width and dimension of the spatial features. Since dense frames are crucial for fine-grained temporal grounding, an appropriate pooling strategy is required to reduce token length. As indicated by Xu et al. (2024a) and Yao et al. (2024) that a 2D average pooling is both efficient and robust for spatial downsampling, we employ a 2D pooling kernel with a size $\sigma_{S}\times\sigma_{S}$ over the feature map and gets $\textbf{F}_{S}\in\mathbb{R}^{N_{S}\times D_{S}}$ as the feature for spatial stream, where $N_{S}=\frac{H_{S}}{\sigma_{S}}\times\frac{W_{S}}{\sigma_{S}}$.
@@
Temporal Stream. Traditional two-stream networks typically encode the optical flow as the temporal stream. However, given the scale of data and parameters of MLLMs, extracting optical flow is computationally expensive and impractical. Consequently, we resort to a strong and well pre-trained video encoder to extract motion representations for each segment, using a lower resolution but more frames. We input each segment, containing $\frac{T}{K}$ frames, into the video encoder to obtain the segment-level features $\textbf{F}_{T}\in\mathbb{R}^{\frac{T}{K}\times H_{T}\times W_{T}\times D_{T}}$, where $H_{T}$, $W_{T}$, $D_{T}$ denote the height, width and dimension of each frame feature. Similar to the spatial stream, we apply a 2D average pooling strategy to downsample $\textbf{F}_{T}$. However, as the temporal stream focuses primarily on temporal modeling, we retain the complete temporal information by only pooling along the spatial dimensions. Specifically, we aggressively downsample $\textbf{F}_{T}$ using a kernel with a larger size of $\sigma_{T}\times\sigma_{T}$, resulting in the compressed $\textbf{F}_{T}\in\mathbb{R}^{\frac{T}{K}\times N_{T}\times D_{T}}$ for temporal stream, where $N_{T}=\frac{H_{T}}{\sigma_{T}}\times\frac{W_{T}}{\sigma_{T}}$. To get the complete segment-level representation, we flatten the features of both stream and concat them together :
$\textbf{F}_{Seg}=\operatorname{Concat}\left[f(\textbf{F}_{S});g(\textbf{F}_{T})\right]$ (1)
@@
where $\textbf{F}_{Seg}\in\mathbb{R}^{(N_{S}+\frac{T}{K}\cdot N_{T})\times D}$, $f(\cdot)$ and $g(\cdot)$ are two MLPs that project the visual features to LLM’s dimension $D$. The final video representation is formed by concatenating the $K$ segment-level representations $\textbf{F}_{Seg}$, resulting in $\textbf{F}_{Vid}\in\mathbb{R}^{K\cdot(N_{S}+\frac{T}{K}\cdot N_{T})\times D}$. This representation retains detailed spatial information across all segments along with their global temporal contexts, while maintaining a manageable token length. The combined video representation $\textbf{F}_{Vid}$ is then fed into the LLM serving as soft prompts, alongside the word embeddings of the instruction text $\textbf{F}_{Text}$ to generate the target response $\mathcal{A}$. The model is trained using the standard cross-entropy loss function for LLM with trainable parameters $\theta$: